# Investigation of the Trailing-Edge Jet-Buffet Phenomenon Using Circulation Control at Transonic Speeds

### v1

@@ Abstract
The phenomenon of a periodic jet detachment occurring on an elliptic airfoil utilizing circulation control at transonic speeds was evaluated using numerical simulations. As the momentum of the jet increases beyond a prescribed magnitude, the jet begins to detach periodically from the surface of the trailing-edge. The effect was analyzed by both URANS calculations as well as by data-driven methods such as the Dynamic Mode Decomposition (DMD). The results of the investigation showed that the effect, termed jet-buffet is decoupled from the dominant shockwave located at the top surface of the wing. Indicating that the jet can no longer augment the wing’s circulation, and thus the termination of circulation control. Furthermore, it was found that the driving mechanism of the jet-buffet originated from a feedback of pressure propagating between the trailing-edge shockwave and the downstream separation bubble produced by the jet. Consequently the jet-buffet causes noticeable fluctuations in the aerodynamic coefficients and such impacts the design implications of circulation-control at transonic flight.
@@ 1 Introduction
Currently, transonic CC numerical research primarily relies on steady jet blowing. For a steady jet, its momentum, expressed as the momentum coefficient must be high enough to prevent premature jet separation. However, detachment behavior is not always consistent across different designs, particularly when using an elliptic trailing-edge, shown in early work by Englar [24] to be favorable for transonic speeds. Experimental results by Alexander [15] revealed that in some cases, jet separation is not immediate but is preceded by a lift plateau. Numerical validations of this behavior [21, 19] indicate that attempting to maintain a steady jet within this plateau region leads to unsteadiness in both residuals and aerodynamic coefficients.
@@
This study will show that the observed lift plateau is a consequence of periodic jet separation and reattachment. This periodic behavior can be numerically captured and analyzed using advanced data-driven methods. Proper Orthogonal Decomposition (POD) [25] enables the decomposition of the flow-field into its dominant coherent structures, while Dynamic Mode Decomposition (DMD) [26] projects these structures onto linearized dynamic modes, facilitating a deeper understanding of unsteady dynamics. These techniques have been extensively used to analyze periodic motion in transonic flows [27, 28, 29], allowing for the identification of key flow features and a more comprehensive study of periodic detachment behavior.
@@
DMD itself is derived by projecting POD modes, obtained through Singular Value Decomposition (SVD), onto the Fourier space. This technique effectively identifies low-rank structures within the data [25]. Since DMD is purely data-driven, it does not require prior knowledge of system dynamics but relies on high-resolution, deterministic data, a requirement linked to its foundation in Koopman operator theory [30]. This constraint on time resolution can, however, be relaxed through sparsity-promoting methods such as compressive sensing [31].
@@
Since unsteady periodic motion can induce responses in aerodynamic forces as well as structural fatigue effects via aeroelastic coupling [32], understanding the periodic detachment can impact transonic CC design methodologies. Thus, this study aims to introduce and characterize the phenomenon, termed jet-buffet, elaborate on its underlying mechanisms, and assess its implications for circulation control.
@@
The paper is structured as follows: Section 2 details the geometry of the model used for calculations, followed by the numerical setup in Section 3. The results in Section 4 present both the steady validation and the characterization of jet-buffet, including its analysis using DMD. Finally, the conclusions of the study are presented in Section 5.
@@ 2 Geometric Model
Figure 1: Geometry of the modeled test wing
The wing cross-section used in this study, depicted in Fig. 1, is based on the geometry employed in circulation control experiments conducted at NASA’s Transonic Dynamic Tunnel (TDT) [15]. It features a 6% thick elliptic airfoil with $0.75\%$ camber and a nominal chord length of $c=30$ inches which was used for computing aerodynamic coefficients. The actual chord length, however, was $28.5$ inches due to a modification where the trailing-edge was cut at $0.9c$ and replaced with a dedicated Coandǎ surface. This replacement surface had an elliptic shape with a major-to-minor axis ratio of $2.98$. The internal plenum, integrated within the wing, was designed to generate the jet required for circulation control. It featured a throat height of $0.012c$ and a length sufficient to produce a fully developed turbulent jet.
The test wing spanned $60$ inches, corresponding to an aspect ratio of $2$, and was secured in place using a splitter-plate reinforced to the tunnel wall. To mitigate three-dimensional effects, an end-plate with a diameter of one chord in length was initially attached to the wingtip. However, numerical validations [19, 33, 21] conducted for transonic circulation control research, along with the author’s prior comparisons with preliminary calculations of the two-dimensional geometry, revealed that both the end-plate and splitter-plate significantly contributed to 3D effects. These effects were observed to influence the strength and position of the upper surface shock at the free-stream Mach number of $M_{\infty}=0.8$. To account for these influences, both the end-plate and splitter-plate were explicitly modeled and their dimensions were selected based on previous numerical studies [21], with values of $1.1c$ and $6c$, respectively. The computational domain was represented as a cylindrical volume with a diameter of $10c$ extending $7c$ in the spanwise direction. To ensure accuracy, the numerical study was validated against the earlier stated experiment at the TDT, which was selected due to its comprehensive dataset covering a wide range of Mach numbers $M$, angles of attack $\alpha$, and jet momentum inputs.
3 Numerical Verification
@@ 3.1 Computational Grid
The computational domain was constructed using an unstructured mesh generated with Fidelity© Pointwise, employing the Voxel algorithm. This resulted in a hybrid grid comprising hexahedral and tetrahedral cells, as illustrated in Fig. 2, which depicts some of the grid’s cross-sections and associated boundary conditions. To study the effects of the grid while accurately capturing the underlying flow physics, an initial fine mesh consisting of $40\times 10^{6}$ cells was utilized, and subsequently coarsened up to $13\times 10^{6}$ cells. Across all grid configurations, the boundary layer was resolved using 80 layers growing at a rate of 1.1, while ensuring the first cell height corresponded to $y^{+}\approx 1$. Four grid variants were developed, focusing on enhanced cell density near the wing, splitter-plate, end-plate, and near-wake regions. For each grid, the mid-span lift coefficient and drag coefficient were computed and compared. Grid convergence was deemed achieved when the variation in these aerodynamic coefficients was within $\mathcal{O}\leq 1\%$. The final grid, consisting of $28\times 10^{6}$ cells, was used for the remainder of the study.
@@ 3.2 Numerical Setup
The computational study was conducted using Ansys© Fluent simulation software by solving the Reynolds Averaged Navier-Stokes (RANS) equations. Boundary conditions were set to replicate the experimental tunnel conditions from  [15] with free-stream Mach number of $M_{\infty}=0.8$, static pressure $p_{\infty}=13789.5\;\mathrm{(Pa)}$ and temperature $T_{\infty}=308.15\;\mathrm{(K)}$ corresponding to a free-stream chord-based Reynolds number of $3\times 10^{6}$. The jet flow from the plenum was modeled as a mass-flow inlet boundary condition, determined by the required momentum coefficient $C_{\mu}=\dot{m}\nu_{j}/{Q_{\mathrm{dyn}}\,c}$, where $\dot{m}$ is the mass-flow to the plenum, $Q_{\mathrm{dyn}}$ is the free-stream dynamic pressure and $\nu_{j}$ is the isentropic jet velocity calculated by the expansion of the jet to the free-stream and defined as:
$\nu_{j}=\sqrt{2R_{g}T_{0,p}\left(\frac{\gamma}{\gamma-1}\right)\left(1-\left(\frac{p_{\infty}}{p_{0,p}}\right)^{\frac{\gamma-1}{\gamma}}\right)}$ (1)
@@
For the numerical schemes, the inviscid fluxes were discretized using 2nd Roe flux difference splitting scheme and the solution gradients were computed using the Green-Gauss node based approach. Both the 2nd order Spalart-Allmaras (SA) and $k-\omega$ SST turbulence models were tested, as depicted in Fig. 3, where the pressure coefficient distribution along the mid-span chord is outlined for both turbulence models and validated against the experimental data, for the cases with no ($C_{\mu}=0$), and with blowing ($C_{\mu}=0.008$).
(a) $C_{\mu}=0$
(b) $C_{\mu}=0.008$
@@
Figure 3: Pressure coefficient distribution at mid-span for both the Spalart-Allmaras and $k-\omega$ SST models, validated against the experimental data at a free-stream Mach number of 0.8 and angle of attack of $3^{\circ}$.
Where: (   experimental results; –.–  SA; –.–  SST)
From a global perspective, CC affects the pressure profile by thickening it and shifting the upper surface shockwave further aft, as shown in Fig. 4 for the Mach field with the SA turbulence model. However, discrepancies were observed between numerical predictions and experimental data regarding the shockwave position for both turbulence models, particularly in the $C_{\mu}=0.008$ case. And it is noted that the $k-\omega$ SST model showed a slightly smaller deviation. It has been suggested that this misalignment is likely due to limitations inherent to the RANS turbulence framework  [21]. Since the experimental setup produced a turbulent boundary layer via an epoxy based trip strip located at 5%c from the leading-edge which was not geometrically modeled.
(a) $C_{\mu}=0$
(b) $C_{\mu}=0.002$
(c) $C_{\mu}=0.004$
(d) $C_{\mu}=0.008$
Figure 4: Mach field at mid-span using the Spalart-Allmaras model, at a free-stream Mach number of 0.8 and angle of attack of $3^{\circ}$.
Focusing on the trailing-edge region, the calculations captured the primary effects of CC on the pressure coefficient distribution as illustrated in Fig. 5 depicting both the pressure distribution along the trailing-edge in Fig. 5(a) as well as its corresponding Mach field in Fig. 5(b). With the addition of blowing, the pressure distribution on the trailing-edge is modified by an addition of two distinct peaks (highlighted by a black box in Fig. 5(b)), which are a consequence of the high velocity jet. This increase in pressure is associated with the formation of a separation bubble. The first peak, located near the plenum outlet (located at $0.95\geq x/c\leq 0.96$) showed a sharper drop than the experimental data. Although no visual data from the report are available, a sharper drop indicates a shorter bubble and thus a misrepresentation by numerical results of the bubble’s length. In contrast, the second peak (located at $0.965\geq x/c\leq 0.99$) was more accurately captured by the SA model. Since the $k-\omega$ SST overly smoothed the solution. Consequently, the Spalart-Allmaras model was selected for the remainder of the study due to its performance in accurately resolving the trailing-edge flow features.
(a) $C_{p}$ distribution
(b) SA based Mach field of the trailing-edge
Figure 5: Validation of the turbulence models at the trailing-edge at Mach number of 0.8, angle of attack of $3^{\circ}$ and momentum coefficient of 0.008.
Where: (   experimental results; –.–  SA; –.–  SST)

### v3

# Analysis of Circulation Control Jet Bi-Stability on a Wing Section at Transonic Speeds via Dynamic Mode Decomposition

@@ Abstract
The phenomenon of stable lift oscillations occurring on an elliptic wing section utilizing circulation control at transonic speeds was evaluated using numerical simulations. As the momentum of the jet increases beyond a prescribed magnitude, periodic detachment occurs from the trailing-edge. This behavior conforms to a bi-stable state, consistent with prior experimental observations. Analysis by both steady and unsteady Reynolds Averaged Navier-Stokes calculations showed that the effect is decoupled from the dominant upstream shockwave. This indicates that the jet can no longer augment the wing’s circulation, marking the termination of circulation control. Furthermore, the results confirm that the absence of the downstream separation bubble acts as the catalyst for this detachment. Dynamic Mode Decomposition analysis revealed that the bi-stability is driven by a pressure feedback between the trailing-edge shockwave and a downstream pressure bubble. A secondary feedback governs the pressure redistribution during the detachment cycle. It was concluded that the pressure-dominant nature of the bi-stability allows it to be captured using relatively simple methods such as URANS, and even approximated through a Reduced Order Model comprising only 2% of the total modes, encapsulating 25% of the modal influence and reconstructing the pressure field with 98% accuracy.
Nomenclature
AA  = time propagation matrix
$\tilde{A}$  = POD projected time propagation matrix
$b$  = DMD initial conditions vector ,(Pa)
$c$  = airfoil chord length, (m)
$c_{d}$  = drag coefficient
$c_{l}$  = lift coefficient
$C_{p}$  = pressure coefficient
$\delta C_{p}$  = pressure coefficient reconstruction error
$C_{\mu}$  = jet momentum coefficient
$\Delta c_{l}$  = incremental lift coefficient due to jet blowing
$f$  = harmonic frequency, (Hz)
$f_{n}$  = natural frequency, (Hz)
II  = modal influence
$I_{\mathrm{id}}$  = identity matrix
$\dot{m}$  = mass flow rate, (kg/s)
$M$  = Mach number
$p$  = pressure, (Pa)
$Q_{dyn}$  = dynamic pressure, (Pa)
$r$  = mode number
$r_{e}$  = trailing-edge major axis, (m)
$Re_{c}$  = chord-based Reynolds number
$R_{g}$  = specific gas constant, (J/(kgK))
$t$  = time ,(s)
$t_{c}$  = bi-stability period, (s)
$t_{c,0}$  = initial detachment time, (s)
$\Delta t$  = time step,(s)
$T$  = temperature ,(K)
$V$  = right singular matrix
$W$  = POD projected eigenvector matrix
$y^{+}$  = non-dimensional wall distance
$x$  = chordwise coordinate, (m)
$X$  = data matrix
$\alpha$  = angle of attack, (deg)
$\gamma$  = specific heat ratio
$\zeta$  = modal damping ratio
$\kappa$  = reduced frequency
$\lambda$  = continuous eigenvalue
$\mu$  = discrete eigenvalue
$\nu_{j}$  = jet isentropic velocity, (m/s)
$\xi$  = exponential decay
$\rho$  = density ,(kg/m3)
$\Sigma$  = singular value matrix
$\tau$  = normalized cycle time
$\Phi$  = modal spatial distribution
$\Psi$  = left singular matrix
Subscripts
0  = total conditions
$p$  = plenum conditions
$\infty$  = freestream conditions
$\mathrm{jet}$  = jet specific properties
@@ 1 Introduction
Currently, transonic CC numerical research primarily relies on steady jet blowing. For a steady jet, its momentum, expressed as the momentum coefficient must be low enough to prevent premature jet separation, limiting its functionality. However, detachment behavior is not always consistent across different designs, particularly when using an elliptic trailing-edge, shown in early work by Englar [24] to be favorable for transonic speeds. Experimental results by Alexander [15] revealed that in some cases, jet separation is not immediate but is preceded by a lift plateau. Numerical validations of this behavior [21, 19] indicate that attempting to maintain a steady jet within this plateau region leads to unsteadiness in both residuals and aerodynamic coefficients.
@@
This study will show that the observed lift plateau is a consequence of periodic jet separation and reattachment. This bi-stability has been depicted previously in experiments regarding the use supersonic jets in quiescent air [25, 26] and thus is an established phenomenon. Nevertheless, as verified throughout this paper, the current bi-stability is heavily influenced by the presence of the transonic free-stream and its consequent trailing-edge shockwave. Furthermore, this periodic behavior can be numerically captured by low-resolution methods such as URANS and analyzed using advanced data-driven methods based solely on its pressure field. Using methods such as Proper Orthogonal Decomposition (POD) [27] enables the decomposition of the flow-field into its dominant coherent structures, while Dynamic Mode Decomposition (DMD) [28] projects these structures onto linearized dynamic modes, facilitating a deeper understanding of unsteady dynamics. These techniques have been extensively used to analyze periodic motion in transonic flows [29, 30, 31], allowing for the identification of key flow features and a more comprehensive study of the periodic detachment behavior.
@@
DMD itself is derived by the eigen-decomposition of the system’s dynamic operator. In practice however, the full decomposition is circumvented by projecting the operator unto the POD modes, obtained through Singular Value Decomposition (SVD), resulting in complex eigenvalues. Such values encode frequency, amplitude and decay rate information, in a similar manner to the classical Fourier transform. This technique effectively identifies low-rank structures within the data [27]. Since DMD is purely data-driven, it does not require prior knowledge of system dynamics but relies on high-resolution, deterministic data, a requirement linked to its foundation in Koopman operator theory [32]. This constraint on time resolution can, however, be relaxed through sparsity-promoting methods such as compressive sensing [33]. The applications of DMD to both analyze the system dynamics as well as to produce a Reduced Order Model (ROM), exemplifies the robustness of a general data-driven approach to system analysis.
@@
Since unsteady periodic motion can induce responses in aerodynamic forces as well as structural fatigue effects via aeroelastic coupling [34], understanding the periodic detachment can impact transonic CC design methodologies. Thus, this study aims to validate the bi-stable phenomenon, elaborate on its underlying mechanisms, analyze whether an ROM can be used to approximate it, and assess its implications for circulation control.
@@
The paper is structured as follows: Section 2 details the geometry of the model used for calculations, followed by the numerical verification and validation of both the steady and unsteady behavior in Section 3. The results in Section 4 present the characterization of bi-stability using URANS, including its analysis, decomposition and reconstruction using DMD. Finally, the conclusions of the study are presented in Section 5.
@@ 2 Geometric Model
Figure 1: Geometry of the modeled TDT experimental wing.
The wing cross-section used in this study, depicted in Fig. 1, is based on the geometry employed in circulation control experiments conducted at NASA’s Transonic Dynamic Tunnel (TDT) [15] and is employed as the validation case for the numerical calculations. The results obtained from the tunnel tests are used, and defined throughout the study as the "TDT tunnel data". The test wing featured a 6% thick elliptic airfoil with $0.75\%$ camber and a nominal chord length of $c=30$ inches which was used for computing aerodynamic coefficients. The actual chord length, however, was $28.36$ inches due to a modification where the trailing-edge was cut at $0.9c$ and replaced with a dedicated Coandǎ surface. This replacement surface had an elliptic shape with a major-to-minor axis ratio of $2.98$, featuring a major axis radius $r_{e}$, of 1.37 inches. The internal plenum, integrated within the wing, was designed to generate the jet required for circulation control. It featured a throat height of $0.012c$ and a length sufficient to produce a fully developed turbulent jet.
The test wing spanned $60$ inches, corresponding to an aspect ratio of $2$, and was secured in place using a splitter-plate reinforced to the tunnel wall. To mitigate three-dimensional effects, an end-plate with a diameter of one chord in length was initially attached to the wingtip. However, numerical validations [19, 35, 21] conducted for transonic circulation control research, along with the author’s prior comparisons with preliminary calculations of the two-dimensional geometry, revealed that both the end-plate and splitter-plate significantly contributed to 3D effects. These effects were observed to influence the strength and position of the upper surface shock at the free-stream Mach number of 0.8. To account for these influences, both the end-plate and splitter-plate were explicitly modeled and their dimensions were selected based on previous numerical studies [21], with values of $1.1c$ and $6c$, respectively. The computational domain was represented as a cylindrical volume with a diameter of $10c$ extending $7c$ in the spanwise direction. To ensure accuracy, the numerical study was validated against the earlier stated experiment at the TDT, which was selected due to its comprehensive dataset covering a wide range of Mach numbers $M$, angles of attack $\alpha$, and jet momentum inputs.
3 Numerical Verification & Validation
@@ 3.1 Computational Grid
The computational domain was constructed using an unstructured mesh generated with Fidelity© Pointwise, employing the Voxel algorithm. This resulted in a hybrid grid comprising hexahedral and tetrahedral cells, as illustrated in Fig. 2, which depicts some of the grid’s cross-sections and associated boundary conditions. To study the effects of the grid while accurately capturing the underlying flow physics, an initial fine mesh consisting of $40\times 10^{6}$ cells was utilized, and subsequently coarsened up to $13\times 10^{6}$ cells. Across all grid configurations, the boundary layer was resolved using 80 layers growing at a rate of 1.1, while ensuring the first cell height corresponded to $y^{+}\approx 1$. Four grid variants were developed, focusing on enhanced cell density near the wing, splitter-plate, end-plate, and near-wake regions. For each grid, the mid-span lift coefficient and drag coefficient were computed and compared as depicted in Fig. 3. Grid convergence was deemed achieved when the variation in these aerodynamic coefficients was within $\mathcal{O}\leq 1\%$. The final grid, consisting of $27\times 10^{6}$ cells, was used for the remainder of the study.
(a) lift grid convergence
(b) drag grid convergence
Figure 3: Convergence of the mid-span lift and drag coefficients with increasing grid count for different mass flow inputs at an angle of attack of 3 and a free-stream Mach number of 0.8.
$N=1.3\times 10^{6}$, $N=1.9\times 10^{6}$, $N=2.7\times 10^{6}$, $N=4\times 10^{6}$
@@ 3.2 Numerical Setup
The computational study was conducted using Ansys© Fluent simulation software by solving the compressible Reynolds Averaged Navier-Stokes (RANS) equations. Boundary conditions were set to replicate the experimental tunnel conditions [15] with free-stream Mach number of $M_{\infty}=0.8$, static pressure $p_{\infty}=13789.5\;\mathrm{(Pa)}$ and temperature $T_{\infty}=308.15\;\mathrm{(K)}$ corresponding to a free-stream chord-based Reynolds number of $2.88\times 10^{6}$. The jet flow from the plenum was modeled as a mass-flow inlet boundary condition, determined by the required momentum coefficient $C_{\mu}=\dot{m}\nu_{\mathrm{jet}}/{Q_{\mathrm{dyn}}\,c}$, where $\dot{m}$ is the mass-flow to the plenum, $Q_{\mathrm{dyn}}$ is the free-stream dynamic pressure and $\nu_{\mathrm{jet}}$ is the isentropic jet velocity calculated by the expansion of the jet to the free-stream and defined as:
$\nu_{\mathrm{jet}}=\sqrt{2R_{g}T_{0,p}\left(\frac{\gamma}{\gamma-1}\right)\left(1-\left(\frac{p_{\infty}}{p_{0,p}}\right)^{\frac{\gamma-1}{\gamma}}\right)}$ (1)
@@
$c_{l}=\frac{1}{c}\int_{0}^{c}(C_{p,\mathrm{low}}-C_{p,\mathrm{up}})dx$ (2)
For the numerical schemes, the inviscid fluxes were discretized using second order Roe flux difference splitting scheme and the solution gradients were computed using the Green-Gauss node based approach. Both the second order Spalart-Allmaras (SA) and $k-\omega$ SST turbulence models were tested to reduce uncertainty and their validation is expanded upon in the following section.
3.3 Turbulence Model & Pressure Validation
(a) no jet blowing, $C_{\mu}=0$
(b) jet blowing with $C_{\mu}=0.008$
Figure 5: Pressure coefficient distribution at mid-span, validated against the TDT experimental data at a free-stream Mach number of 0.8 and angle of attack of $3^{\circ}$, as well as the respective Mach fields using the SA turbulence model.
Spalart-Allmaras,   $k-\omega$ SST,   Chen et al. [23],   TDT tunnel data
Application of circulation control augments the pressure around the airfoil surface, resulting in a wider pressure profile and a shift in the upper surface shockwave further aft, as shown in Fig. 5 which depicts the pressure coefficient distribution for both no blowing ($C_{\mu}=0$) and with jet blowing ($C_{\mu}=0.008$) as well as the Mach fields of each via the SA model. The calculations reproduce the test data sufficiently across the airfoil surface and exceptionally at the trailing-edge region, yet a discrepancy was observed between numerical predictions and experimental data regarding the shockwave position for both turbulence models. This misalignment is observed particularly using $C_{\mu}=0.008$, and it is noted that the $k-\omega$ SST model showed a slightly smaller deviation.
The misalignment has been encountered in similar studies [21, 18, 19], and the results of the current simulations and those of Chen et al [23] are compared to illustrate this inconsistency across different calculations. Forester et al [21] suggested that its presence could be due to limitations in assuming a fully turbulent boundary layer, since the experimental setup produced a turbulent boundary layer via an epoxy based trip strip located at 5%c from the leading-edge which was not geometrically modeled. This conclusion however, is not consistently reflected across prior work. With no uncertainty analysis present in the original test report, an equivalent case could be made for inaccuracies due to sensor limitations within the sensitive shock bubble. Nevertheless, offering predominantly accurate predictions, the current study is believed to reproduce the pressure field with sufficient fidelity for further analysis.
3.4 Steady State Validation
Figure 6: Effect of momentum addition on the mid-span lift enhancement at a free-stream Mach number of 0.8 and an angle of attack of $3^{\circ}$ and $0^{\circ}$.
@@ 3.4 Steady State Validation
From an integral perspective, the lift enhancement, $\Delta c_{l}$ is an important parameter, as it isolates the nominal lift of the wing from the lift gained via CC, dictating the efficacy of the of the jet injection on the external aerodynamics. Validation of this feature is presented in Fig. 6 which illustrates the variation of the lift enhancement as functions of the momentum coefficient at an angles of attack of $3^{\circ}$ and $0^{\circ}$ and a freestream Mach number of $0.8$.
Accurate prediction is achieved compared with the TDT data, especially in the linear region region of the momentum coefficient ($0\leq C_{\mu}\leq 0.006$) with a root mean square error (RMSE) of 1.21% and 2.75% for $\alpha=0^{\circ},3^{\circ}$ respectively. The momentum coefficient value of 0.008 was the largest validated in steady-state with a relative error of 7%, as the augmentation trend trend shifts from linear to a an almost constant behavior with increasing jet strength.
When the momentum input increases beyond a certain threshold ($C_{\mu}=0.0137$), the jet loses its authority, detaches from the surface, and the lift enhancement rapidly declines towards zero with increasing jet strength, preceded by a plateau. In the plateau region (at momentum inputs of $C_{\mu}>0.008$), attempts to achieve a steady-state solution revealed unsteady behavior in both the aerodynamic coefficients and the residuals. Similar observations were made in other studies [21, 19], although no detailed explanation was provided for this phenomenon.
3.5 Unsteady Validation
To evaluate the unsteadiness at the plateau, second order Unsteady Reynolds-Averaged Navier-Stokes (URANS) simulations were performed at an angle of attack of $3^{\circ}$ and a momentum coefficient of 0.01, corresponding to the first indication of unsteady behavior. The results were verified through a time step sensitivity study, which considered three different time increments: $\Delta t=(1\times 10^{-2},5\times 10^{-3},1\times 10^{-3})\;\mathrm{(ms)}$, equating to $(128,256,1500)$ iterations per cycle, respectively. Convergence was assessed based on a three-order-of-magnitude reduction in residuals per time step, as well as the stabilization of the lift coefficient’s amplitude and frequency, determined via a one-sided Fast Fourier Transform (FFT). Among the tested time increments, $\Delta t=5\times 10^{-3}\;\mathrm{(ms)}$ was deemed sufficient for convergence, yielding a lift coefficient amplitude of 0.032 and a reduced frequency $\kappa_{\mathrm{jet}}=f_{\mathrm{jet}}r_{e}/\nu_{\mathrm{jet}}$ of 0.105. Here, $f_{\mathrm{jet}}$ denotes the jet periodic detachment’s natural frequency.
Although no transient data from the TDT experiment was available, a quantitative comparison can nonetheless be performed. This is achieved by computing an ensemble average of the CFD solution over one full cycle and comparing it to the frequency-averaged data provided in the report. Fig. 7 presents the time-averaged CFD pressure distribution at $C_{\mu}=0.01$ and the steady value of $C_{\mu}=0.008$ alongside the experimental measurements taken near the trailing-edge.
(a) wing
(b) trailing-edge
@@ 3.5 Unsteady Validation
Figure 7: Pressure coefficient distributions at mid-span, at a free-stream Mach number of 0.8 and angle of attack of $3^{\circ}$ of the time ensemble CFD data and the steady state data as compared with the TDT experimental result.
time ensemble at $C_{\mu}=0.01$,   $C_{\mu}=0.008$,   TDT tunnel data at $C_{\mu}=0.011$
A clear similarity is observed up to the trailing-edge ($0\leq x/c\leq 0.95$) depicted in Fig. 7(a) between the steady-state solution at $C_{\mu}=0.008$ and the time ensemble at $C_{\mu}=0.01$, where the pressure distributions align closely. Notably, in contrast to the behavior at lower jet momentum values ($C_{\mu}\leq 0.008$), increasing the momentum to the onset of aerodynamic instability does not alter the position of the upper surface shock wave ($0\leq x/c\leq 0.4$). This suggests that when ejecting a jet of this value, circulation control can no longer augment the upper surface pressure distribution upstream of the trailing-edge. At the trailing-edge region ($0.95\leq x/c\leq 1$) depicted in Fig. 7(b) , the pressure peak present at ($0.97\leq x/c\leq 0.99$) using $C_{\mu}=0.008$ is absent at $C_{\mu}=0.01$. As the primary feature distinguishing the distributions, its disappearance must pertain to the onset of jet detachment.
(a) jet attachment
(b) jet detachment
Figure 8: Jet bi-stability obtained from the URANS calculations and compared to the similar case of bi-stability at quiescent air [25].
Further analysis of the flow-field behavior near the trailing-edge indicates that the observed unsteadiness results from the periodic detachment of the jet, as shown in Fig. 8 for both the detachment and reattachment phases. The resulting bi-stability closely resembles the well-substantiated periodic separation observed in experimental studies of supersonic jet ejection into quiescent air around a circular trailing-edge [25, 26], also illustrated in Fig. 8. This similarity is further supported by their use of a comparable nozzle pressure ratio of 3.6, relative to 3.96 in the present study. Therefore, it is reasonable to deduce that the effect observed here represents the same form of bi-stability. However, unlike the quiescent air experiments, the current phenomenon is constrained by the presence of the trailing-edge shockwave due to the elliptical geometry and transonic freestream, which terminates the detachment rather than allowing it to persist until natural equilibrium is reached.